show work for each individual resister for a and b

A 12.0 V battery is hooked up to two resistors in series: a 1000 Ohm resistor and a 2200 Ohm resistor.
(a) Calculate the voltage across each resistor.
(b) Calculate the current through each resistor.
(c) What is the total resistance of the circuit?
Answer #1

here,

R1 = 1000 ohm

R2 = 2200 ohm

voltage of battery , V = 12 V

a)

as they are connected in series , the total resistance of circuit , R = R1 + R2

R = 1000 ohm + 2200 ohm

R = 3200 ohm

the current in through each resistor , I = V /R

I = 12 /3200 A = 3.75 * 10^-3 A

the resistance across R1 , V1 = R1 * I

V1 = 3.75 * 10^-3 * 1000 = 3.75 V

the resistance across R2 , V2 = V - V1

V2 = 8.5 V

b)

the current in through each resistor , I = V /R

I = 12 /3200 A = 3.75 * 10^-3 A

c)

What is the total resistance of the circuit is 3200 ohm